OpenStudy (anonymous):

Arc Length? @dan815 @ganeshie8

OpenStudy (anonymous):

Arc Length

OpenStudy (anonymous):

Geometry

OpenStudy (jhannybean):

 \(\large s\) = arc length , \[\large s= \cfrac{\theta}{360} \cdot 2 \pi r\]

OpenStudy (anonymous):

Whats the Radius ?

OpenStudy (jdoe0001):

hehe

OpenStudy (jhannybean):

\(r\) = radius....

OpenStudy (jdoe0001):

you only asked for the Arc length generically

OpenStudy (anonymous):

\[s = 5/360 \]

OpenStudy (jhannybean):

Do you have a specific question perhaps?

OpenStudy (anonymous):

Times 2PIR

OpenStudy (anonymous):

yea

OpenStudy (jhannybean):

How did you find that? You never gave us any of the variables to work with.

OpenStudy (anonymous):

Something like that

OpenStudy (jhannybean):

Oh okay.

OpenStudy (anonymous):

Radius is 12

OpenStudy (jhannybean):

\[\large s = \cfrac{90}{360} \cdot 2( \pi) (12)\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

18.85

OpenStudy (jhannybean):

\[\large s= \cfrac{1}{4} \cdot 24 \pi = 6 \pi \]

OpenStudy (jhannybean):

you're correct :) good job.

OpenStudy (anonymous):

jhanny

OpenStudy (anonymous):

I need guidance to see if my answers are correct -.-

OpenStudy (anonymous):

If I just get 1 wrong my grade goes down -.- on the practices

OpenStudy (anonymous):

Wait is it CM? The units?

OpenStudy (jhannybean):

You've correctly inserted the variables into the right places in the formula, and you solved it correctly where your answer equaled to mine. You're good to go.

OpenStudy (jhannybean):

cm=centimeters.

OpenStudy (anonymous):

yea

OpenStudy (dan815):

flazingblazes do u remember everything we learnt so far

OpenStudy (anonymous):

Yes.

OpenStudy (dan815):

how many degrees in a circle

OpenStudy (anonymous):

Call me Tim

OpenStudy (anonymous):

360

OpenStudy (dan815):

ok

OpenStudy (jhannybean):

Tim.

OpenStudy (dan815):

okay now 360 degrees = 2pi radians

OpenStudy (dan815):

have you learnt about that yet?

OpenStudy (anonymous):

Circumference?